OpenStudy (anonymous):

Now, if u = tan^(−1)2x, then (1/2)int tan^(−1)2x/1 + 4x dx= u du

OpenStudy (anonymous):

ok... something wrong about it... did you mean .../(1+4x^2)...?

OpenStudy (anonymous):

never mind i got it thanks

OpenStudy (anonymous):

I assume that you mean: \[\int\limits_{ }^{}\arctan(2x)/(1+4x^2) dx=1/2\int\limits_{ }^{}udu=1/4 u^2 +const=...\] good Luck!

OpenStudy (anonymous):

this is 1/2arctan(2x)^2+C